OpenStudy (anonymous):

Which of the following expressions would appear furthest to the right on a number line? A,√3+5 B.7 C,√36 D.15/2

OpenStudy (angela210793):

15/2

OpenStudy (anonymous):

15/2 = 7.5

OpenStudy (anonymous):

didn't we see these yesterday exactly?

OpenStudy (angela210793):

sqr3+5=6.7 7 6 7.5 so is 15/2

OpenStudy (anonymous):

i've seen a few questions today that were exactly from yesterday

OpenStudy (across):

\[\sqrt{3}<2\rightarrow2+5<7\]\[7=7\]\[\sqrt{36}=6\]\[15/2=7.5\]7.5 is the greatest.

OpenStudy (anonymous):

so d ?

OpenStudy (across):

yes

OpenStudy (anonymous):

It's probably another round of students doing computer based math programs like the PLATO group we had not too long ago.